# Annotator-Centric Active Learning for Subjective NLP Tasks

### v1

@@ A.4 ACAL Annotator Strategy details
uses a sentence embedding model to represent the content that an annotator has annotated. We use all-MiniLM-L6-v2111https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. We select annotators that have not annotated yet (empty history) before picking from those with a history to prioritize filling the annotation history for each annotator.
@@ A.5 Disagreement rates
We report the average disagreement rates per dataset and task in Figure 5, for each of the dataset and task combinations.
@@ Appendix B Detailed Results Overview
B.1 Test set evaluation other MFTC and MHS tasks
See Table 6 for the trade-off between data efficiency and test-set performance for the two conventional metrics ( and ). We include copy the earlier mentioned results for MFTC (care) and MHS (dehumanize) for convenience.
B.2 Annotator-Centric evaluation for other MFTC and MHS tasks
We show the full annotator-centric metrics results for MFTC betrayal (Table 8), MFTC loyalty (Table 9), MHS genocide (Table 10), and MHS respect (Table 11).
Average | Worst-off | ||||
App. | |||||
0.578 | 0.147 | 0.420 | 0.199 | -1.6 | |
0.581 | 0.149 | 0.433 | 0.212 | -1.6 | |
0.593 | 0.161 | 0.430 | 0.239 | -5.0 | |
0.583 | 0.148 | 0.429 | 0.199 | -1.6 | |
0.594 | 0.150 | 0.419 | 0.203 | -2.5 | |
0.584 | 0.148 | 0.434 | 0.200 | -1.3 | |
0.588 | 0.149 | 0.435 | 0.204 | -1.0 | |
0.591 | 0.149 | 0.428 | 0.194 | -2.6 | |
0.589 | 0.147 | 0.431 | 0.195 | -48.6 | |
0.589 | 0.149 | 0.430 | 0.200 | -0.0 | |
passive | 0.481 | 0.199 | 0.360 | 0.290 | 0.0 |
Average | Worst-off | ||||
App. | |||||
0.564 | 0.177 | 0.222 | 0.372 | -0.4 | |
0.563 | 0.176 | 0.222 | 0.374 | -0.3 | |
0.573 | 0.176 | 0.222 | 0.370 | -0.3 | |
0.551 | 0.175 | 0.222 | 0.373 | -0.3 | |
0.557 | 0.177 | 0.217 | 0.357 | -1.1 | |
0.541 | 0.177 | 0.222 | 0.355 | -0.8 | |
0.556 | 0.177 | 0.222 | 0.358 | -0.9 | |
0.558 | 0.177 | 0.222 | 0.358 | -1.3 | |
0.560 | 0.176 | 0.222 | 0.361 | -29.1 | |
0.559 | 0.177 | 0.222 | 0.366 | -0.1 | |
passive | 0.512 | 0.183 | 0.261 | 0.309 | 0.0 |
Average | Worst-off | ||||
App. | |||||
0.700 | 0.227 | 0.000 | 0.560 | -6.3 | |
0.698 | 0.225 | 0.000 | 0.565 | -1.7 | |
0.700 | 0.224 | 0.000 | 0.566 | -1.7 | |
0.702 | 0.224 | 0.000 | 0.565 | -1.7 | |
0.711 | 0.229 | 0.000 | 0.549 | -12.6 | |
0.707 | 0.231 | 0.000 | 0.548 | -7.9 | |
0.709 | 0.231 | 0.000 | 0.548 | -7.9 | |
0.712 | 0.229 | 0.000 | 0.547 | -12.6 | |
0.339 | 0.387 | 0.000 | 0.496 | -60.1 | |
0.331 | 0.390 | 0.000 | 0.497 | -24.7 | |
passive | 0.700 | 0.245 | 0.000 | 0.570 | – |
Average | Worst-off | ||||
App. | |||||
0.460 | 0.331 | 0.000 | 0.528 | -18.8 | |
0.456 | 0.331 | 0.000 | 0.530 | -18.8 | |
0.461 | 0.331 | 0.000 | 0.529 | -18.8 | |
0.460 | 0.331 | 0.000 | 0.528 | -18.8 | |
0.466 | 0.323 | 0.000 | 0.533 | -4.9 | |
0.463 | 0.323 | 0.000 | 0.532 | -4.9 | |
0.459 | 0.324 | 0.000 | 0.531 | -4.9 | |
0.462 | 0.324 | 0.000 | 0.532 | -4.9 | |
0.339 | 0.387 | 0.000 | 0.496 | -60.1 | |
0.331 | 0.390 | 0.000 | 0.497 | -24.7 | |
passive | 0.259 | 0.405 | 0.000 | 0.587 | – |
B.3 Training process
In our main paper, we report a condensed version of all metrics during the training phase of the active learning approaches. Below, we provide a complete overview of all approaches over all metrics. The results can be seen in Figures 7 through 13.

### v2

@@ A.4 ACAL annotator strategy details
uses a sentence embedding model to represent the content that an annotator has annotated. We use all-MiniLM-L6-v2222https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. We select annotators that have not annotated yet (empty history) before picking from those with a history to prioritize filling the annotation history for each annotator.
@@ A.5 Disagreement rates
We report the average disagreement rates per dataset and task in Figure A1, for each of the dataset and task combinations.
[width=]tikz/entropies
@@ Appendix B Detailed results overview
B.1 Annotator-Centric evaluation for other MFTC and MHS tasks
We show the full annotator-centric metrics results for MFTC betrayal, MFTC loyalty, MHS genocide, and MHS respect in Table B1. This follows the same format at Table 1. The results in this table also form the basis for Figure 5.
Average | Worst-off | |||||||
App. | ||||||||
MFTC (betrayal) | 71.5 | .047 | 57.8 | .147 | 42.0 | .199 | -1.6 | |
71.2 | .046 | 58.1 | .149 | 43.3 | .212 | -1.6 | ||
71.2 | .051 | 59.3 | .161 | 43.0 | .239 | -5.0 | ||
71.0 | .046 | 58.3 | .148 | 42.9 | .199 | -1.6 | ||
72.6 | .042 | 59.4 | .150 | 41.9 | .203 | -2.5 | ||
73.6 | .045 | 58.4 | .148 | 43.4 | .200 | -1.3 | ||
74.0 | .045 | 58.8 | .149 | 43.5 | .204 | -1.0 | ||
73.2 | .044 | 59.1 | .149 | 42.8 | .194 | -2.6 | ||
72.1 | .046 | 58.9 | .147 | 43.1 | .195 | -48.6 | ||
71.8 | .047 | 58.9 | .149 | 43.0 | .200 | -0.0 | ||
PL | 75.2 | .037 | 48.1 | .199 | 36.0 | .290 | 0.0 | |
MFTC (betrayal) | 66.9 | .034 | 56.4 | .177 | 22.2 | .372 | -0.4 | |
68.9 | .032 | 56.3 | .176 | 22.2 | .374 | -0.3 | ||
67.1 | .031 | 57.3 | .176 | 22.2 | .370 | -0.3 | ||
68.4 | .031 | 55.1 | .175 | 22.2 | .373 | -0.3 | ||
61.3 | .032 | 55.7 | .177 | 21.7 | .357 | -1.1 | ||
66.5 | .032 | 54.1 | .177 | 22.2 | .355 | -0.8 | ||
62.4 | .033 | 55.6 | .177 | 22.2 | .358 | -0.9 | ||
64.4 | .031 | 55.8 | .177 | 22.2 | .358 | -1.3 | ||
71.5 | .030 | 56.0 | .176 | 22.2 | .361 | -29.1 | ||
66.5 | .033 | 55.9 | .177 | 22.2 | .366 | -0.1 | ||
PL | 62.5 | .029 | 51.2 | .183 | 26.1 | .309 | 0.0 | |
MHS (genocide) | 26.5 | .050 | 70.0 | .227 | 0.0 | .560 | -6.3 | |
28.2 | .051 | 69.8 | .225 | 0.0 | .565 | -1.7 | ||
28.1 | .051 | 70.0 | .224 | 0.0 | .566 | -1.7 | ||
28.3 | .050 | 70.2 | .224 | 0.0 | .565 | -1.7 | ||
32.8 | .077 | 71.1 | .229 | 0.0 | .549 | -12.6 | ||
27.7 | .048 | 70.7 | .231 | 0.0 | .548 | -7.9 | ||
26.7 | .048 | 70.9 | .231 | 0.0 | .548 | -7.9 | ||
27.3 | .048 | 71.2 | .229 | 0.0 | .547 | -12.6 | ||
28.0 | .048 | 33.9 | .387 | 0.0 | .496 | -60.1 | ||
33.3 | .080 | 33.1 | .390 | 0.0 | .497 | -24.7 | ||
PL | 21.6 | .044 | 70.0 | .245 | 0.0 | .570 | – | |
MHS (respect) | 41.4 | .086 | 46.0 | .331 | 0.0 | .528 | -18.8 | |
40.8 | .087 | 45.6 | .331 | 0.0 | .530 | -18.8 | ||
41.2 | .086 | 46.1 | .331 | 0.0 | .529 | -18.8 | ||
40.6 | .086 | 46.0 | .331 | 0.0 | .528 | -18.8 | ||
32.8 | .077 | 46.6 | .323 | 0.0 | .533 | -4.9 | ||
41.0 | .085 | 46.3 | .323 | 0.0 | .532 | -4.9 | ||
41.8 | .084 | 45.9 | .324 | 0.0 | .531 | -4.9 | ||
40.6 | .085 | 46.2 | .324 | 0.0 | .532 | -4.9 | ||
41.7 | .085 | 33.9 | .387 | 0.0 | .496 | -60.1 | ||
33.3 | .080 | 33.1 | .390 | 0.0 | .497 | -24.7 | ||
PL | 41.0 | .080 | 25.9 | .405 | 0.0 | .587 | – |
B.2 Training process
In our main paper, we report a condensed version of all metrics during the training phase of the active learning approaches. Below, we provide a complete overview of all approaches over all metrics. The results can be seen in Figures B1 through B7.
[width=]tikz/training_process_DICES_overall
[width=]tikz/training_process_MFTC_care
[width=]tikz/training_process_MFTC_loyalty
[width=]tikz/training_process_MFTC_betrayal
[width=]tikz/training_process_MHS_dehumanize
[width=]tikz/training_process_MHS_genocide
[width=]tikz/training_process_MHS_respect